OpenStudy (anonymous):

How do I find the domain of the function in this problem? f(x)=x/3x-1

OpenStudy (anonymous):

The domain required, if we're talking about only real numbers here, is when imputing an x into the denominator to make the denominator NOT 0. Therefore, the domain is when \[x \neq 1/3\]

OpenStudy (anonymous):

Now what about the infinite union?

OpenStudy (mathteacher1729):

Now what about the infinite union? If your function is \[ f(x)=\frac{x}{3x-1} \] the domain is exactly \[x\neq \frac{1}{3}\]. Representing this as a union would be: \[(-\infty , \frac{1}{3}) \cup (\frac{1}{3}, \infty \]

OpenStudy (mathteacher1729):

I forgot to close the parenthesis: \[(-\infty, \frac{1}{3}) \cup (\frac{1}{3}, \infty)\]